OpenStudy (anonymous):

a particle moves along the y axis with a motion defined by the equation s(t)= lnt/t for t>0 find the velocity at t=2 PLEASE HELP!

OpenStudy (anonymous):

v=ds/dt=[1-ln t]/t^2 v(2)=[1-ln2]/4 ans v(2)=(1-0.693)/4 =0.0767 ans